# Groups with et0l co-word problem

### v1

@@ 6 Closure properties of special co-et0l-groups
Otherwise choose $i$ such that the component $w_{i}$ of $w$ lies within the word problem of one of $H$ or $K$, and define $u:=w_{1}\cdots w_{i-1}$, $u^{\prime}:=w_{i+1}\cdots w_{k}$, $v:=w_{i}$; if $i=1$ or $i=k$ then $u$ or $u^{\prime}$ is defined to be empty. Since, by assumption, we have $w=uvu^{\prime}\in\mathsf{coWP}(G,A)$ and $v=_{G}1$, we have $uu^{\prime}\in\mathsf{coWP}(G,A)$ and $v\in\mathsf{WP}(G,A)$. By inductive hypothesis, $uu^{\prime}$, which has at most $k-1$ components, is within $\mathcal{L}(\mathcal{M}_{G})$, so there is an entry configuration $C$ of $\mathcal{M}_{G}$ for which $C^{uu^{\prime}}$ is an accept configuration. Furthermore, for any $n>0$, we can choose $C$ to satisfy Condition (ii) above, and in any case we choose it to satisfy this condition with $n$ equal to the length of $w$.
@@
Now, since $\mathcal{M}_{G}$ ignores the word problem (i.e. we have Property (3)) and $C^{uv}$ does not fail, we have $C^{w}=C^{uvu^{\prime}}=C^{uu^{\prime}}$ is an accept state, and $C$ satisfies Condition (ii) above as required. ∎

### v2

@@ 6 Closure properties of special co-et0l groups
Otherwise choose $i$ such that the component $w_{i}$ of $w$ lies within the word problem of one of $H$ or $K$, and define $u\coloneqq w_{1}\cdots w_{i-1}$, $u^{\prime}\coloneqq w_{i+1}\cdots w_{k}$, and $v\coloneqq w_{i}$; if $i=1$ or $i=k$ then $u$ or $u^{\prime}$ is defined to be empty. Since, by assumption, we have $w=uvu^{\prime}\in\mathsf{coWP}(G,A)$ and $v=_{G}1$, we have $uu^{\prime}\in\mathsf{coWP}(G,A)$ and $v\in\mathsf{WP}(G,A)$. By inductive hypothesis, $uu^{\prime}$, which has at most k1k-1 components, is within $\mathcal{L}(\mathcal{M}_{G})$, so there is an entry configuration $C$ of $\mathcal{M}_{G}$ for which $C^{uu^{\prime}}$ is an accept configuration. Furthermore, for any $n>0$, we can choose $C$ to satisfy Condition (ii) above, and in any case we choose it to satisfy this condition with $n$ equal to the length of $w$.
@@
Now, since $\mathcal{M}_{G}$ ignores the word problem (i.e. we have Property (P6)) and $C^{uv}$ does not fail, we have $C^{w}=C^{uvu^{\prime}}=C^{uu^{\prime}}$ is an accept configuration, and $C$ satisfies Condition (ii) above as required. ∎